# Homomorphism, substructure and ideal: Elementary but rigorous aspects of renormalization group or hierarchical structure of topological orders

### v1

@@ Abstract
We study ring homomorphisms between fusion rings appearing in conformal field theories connected under massless renormalization group (RG) flows. By interpreting the elementary relationship between homomorphism, quotient ring, and projection, we propose a general quantum Hamiltonian formalism of a massless and massive RG flow with an emphasis on generalized symmetry. In our formalism, the noninvertible nature of the ideal of a fusion ring plays a fundamental role as a condensation rule between anyons. Our algebraic method applies to the domain wall problem in $2+1$ dimensional topologically ordered systems and the corresponding classification of $1+1$ dimensional gapped phase, for example. An ideal decomposition of a fusion ring provides a straightforward but strong constraint on the gapped phase with noninvertible symmetry and its symmetry-breaking (or emergent symmetry) patterns. Moreover, even in several specific homomorphisms connected under massless RG flows, less familiar homomorphisms appear, and we conjecture that they correspond to partially solvable models in recent literature. Our work demonstrates the fundamental significance of the abstract algebraic structure, ideal, for the RG in physics.
@@
Classification of the topological orders (TOs) and their fusion rules (or fusion rings[1]) between anyons is one of the most fundamental interests in contemporary condensed matter physics and related research fields[2, 3, 4, 5, 6, 7]. Recently, the transformation law of the anyons in topologically ordered systems through a domain wall, which has been studied in the Kondo or impurity problem[8, 9, 10, 11, 12, 13, 14, 15, 16] traditionally, has captured attention in the fields as a potential classification framework of TOs more commonly (We note [17, 18, 19] as useful references for traditional research directions. We also note older works in astrophysics[20, 21] from a historical view). There exist mutually related ideas to study such settings, such as gapped or symmetry preserving domain walls[22, 23, 24, 25, 26, 27, 28], Witt equivalence[29, 30, 31], and renormalization group (RG) domain walls[32, 33]. More practically, it is known that the theory containing a domain wall can explain the controversy of thermal Hall conductance in experimental settings[34, 35] and this is a consequence of anomaly-inflow mechanism[21, 36, 37, 38]. In other words, if anyons in one theory can smoothly transform into anyons in another theory, these two theories are connected in some sense and a nontrivial domain wall between them can exist. In this sense, bulk edge correspondence[2, 5, 6] or the correspondence between conformal field theory (CFT) and topological quantum field theory (TQFT)[39], CFT/TQFT in short, should be relaxed to this connectivity or equivalence relations when applying the field-theoretic ideas to a realistic system as in experimental settings.
@@
On the other hand, such connectivity between theories is a close connection to symmetry-protected topological order (SPT) [40, 41, 42, 43, 44, 45, 46, 47, 48, 48, 49, 50, 51, 52, 53, 54] and its gapless analog called gapless SPT[55, 56, 57, 58] by the CFT/TQFT correspondence[39]. It has been gradually established that t’Hooft anomaly[59] or Lieb-Shultz-Mattis anomaly[60, 61] classification governs such connectivity between theories[62, 63, 64, 65, 66, 67, 68, 69] and the corresponding phenomena has been studied numerically in [70, 71, 72]. The most fundamental object in this classification is the integer spin simple current proposed by Schellekens and the collaborators[73, 74, 75, 76, 77, 78, 79, 80] which initially appeared in the study of discrete torsion[81, 82]. If the product of two theories forms an integer (or half-integer) spin simple current analogous to hadron, the product theory can be extended or gauged with respect to this nontrivial integer spin simple current. By applying the folding trick[12], the product or coupled theory generates a class of domain walls between the theories phenomenologically. Hence, one may expect a systematic construction of anyon transformation laws between theories in the same anomaly class under such anomaly classification by the integer spin simple current. However, the research has been still limited at this stage (Probably, the fundamental problem in this research direction is the appearance of mixing of chirality of anyons in the coupled theories. To our knowledge, this object has been studied in the context of the fractional quantum Hall effect in higher Landau levels or anti-Pfaffian states [83, 84, 85]. See also the discussions in [86, 69]).
@@
In this work, we demonstrate a general phenomenology realizing an RG as a projection. For the construction of this projection, a mathematical structure called ideal plays a fundamental role. An ideal is fundamentally noninvertible and nonabelian (or nongroup-like). Moreover, it is also different from subring which commonly appears in physics. In principle, one can construct (infinitely) many new transformation laws of anyons, by considering the ideal decomposition of existing fusion rings. More respectively, we study anyon transformation law or ring homomorphism between anyons in conformal field theories connected by massless RG flows[87]. Surprisingly, even in the simplest flow, such as the $SU(2)_{1}\times SU(2)_{1}$ Wess-Zumino-Witten (WZW) model to the $SU(2)_{2}$ WZW model and the flow from the tricritical Ising CFT to the Ising CFT, negative and fractional fusion coefficient, and unusual (or nonintegrable) flows appear inevitably. The unusual fusion coefficient has appeared recently in [88, 89, 58, 90, 91], and the unusual flow seems to correspond to partially solvable systems in [92, 93, 94].
@@
The rest of the manuscript is organized as follows. In section II, we provide a general or mathematical understanding of massless RG flow and its implication to gapped phases in $1+1$ dimensions and that to bulk states in topological orders in $2+1$ dimensions. This is the main section of this work and we demonstrate the fundamental importance of further studies on ideal structure in fusion ring. The argument is purely algebraic and one can generalize the argument to higher dimensional systems. In section III, we introduce a way of constructing a series of ideals from a given ring corresponding to a CFT. The method itself is well known in mathematical literature, but its implication for theoretical physics is quite strong. This construction itself implies the existence of (infinitely) many new series of massless RGs from a given CFT. In section IV, we provide several detailed constructions of homomorphism between fusion rings appearing in CFTs connected under massless RG flows. We demonstrate the existence of nontrivial homomorphisms that are outside of existing integrable massless RG flow and propose its possible connection to partially solvable systems in recent literature[92, 93, 94]. Section. V is the concluding remarks of this manuscript and we remark on a possible extension of our research and related phenomena that will be realizable in experimental settings. In Appendix A, we present an alternative way to obtain a Hamiltonian realizing the massless RG by taking some limit. The resultant model can be interpreted as a gauge model involving the ideal structures. In Appendix B, we provide a conjectural decomposition of CFTs related by massless RG which will be useful in studying the homomorphism between models.
II Substructure from homomorphism and projection as renormalization group flow
@@
where $\mathbf{A}$ is an algebra describing generalized symmetries (or integral of motions) or underlying fusion category of anyons and the lower index $(1)$ and $(2)$ label the ultra-violet (UV) and infra-red (IR) theories respectively. The corresponding argument has appeared evidently in [33] based on the earlier work [96] and folding trick [12]. When interpreting this homomorphism as a defect between two theories, it is called RG domain wall and it is expected to describe a massless RG[87]. We also note that the categorical structure corresponding to the ring homomorphism, called functor, has been studied as gapped (or symmetry-preserving) domain walls in [23, 24, 26, 28] with emphasis on the theoretical aspect of TQFTs or TOs.
@@
We provide an interpretation of the algebraic analog of a coset of $\mathbf{A}_{(1)}$ by $\text{Ker}\rho$ or the quotient ring $\mathbf{A}_{(1)}/\text{Ker}\rho$ which can be constructed from the homomorphism $\rho$ corresponding to the massless RGs. We note that the idea of introducing coset by group symmetry, known as orbifolding has been studied widely[97, 98, 99, 81, 100, 101, 102] and is generalized to nonabelian symmetry[103, 104], but the quotient ring which we will discuss is different from them. The object, $\text{Ker}\rho$, deviding a ring is an ideal and a nontrivial ideal only forms a pseudo ring or rng without identity. (We cite [105] as a recent english translation of the fundamental work by Noether [106] to emphasize the fundamental importance of ideal in contemporary mathematics.) In other words, an ideal is intrinsically noninvertible or nonabelian (or non-group-like). From these properties, it should be distinguished from a subring which is familiar also in physics literature. We also note that the symbol “///" in this section is different from “///" in coset CFTs[107, 108, 109, 110], whereas there exist some connections between the models as in [111, 33] (We introduce a conjectural relationship between coset CFTs and massless RGs in the Appendix. B). As far as we know, the term ideal appeared in physics literature in [112] in the study of matrix product operator, in [113] with a close connection to massless RG flows and in [114] in studying polynomial ring representations of fractonic systems[115, 116].
@@
We restrict our attention to abstract algebra, not to category theory because the underlying category theories can be premodular fusion or superfusion categories or even outside of them that are still under development. However, the basic fusion ring in this manuscript is well-defined and has been studied several times both in mathematics and physics communities. In other words, our discussion can be a clue to constructing a new fundamental class of categories. The fusion rings with negative or fractional coefficients are the corresponding structure and they appeared in the study of fermionic string theories [117] and more recently in [89, 58], for example. The corresponding fusion rings have been studied more generally by the first author in [90, 91] with a close connection to the simple current extension by Schellekens and the collaborators[73, 74, 75, 76, 77, 78, 79, 80].
@@
where $\text{Ker}\rho$, forming an ideal, is a kernel of the map $\rho$ and the inclusion is defined by interpreting the objects as linear algebra (not as a ring). This is widely known as ring homomorphism theorem in mathematics (or an example of Noether’s isomorphism theorems) and one can see the profound history initiated by Kummer, Dedekind, and more. In the present theoretical physics context, this means the correspondence between $\mathbf{A}_{(1)}/\text{Ker}\rho$ and the IR generalized symmetries (or integral of motions) of massless flow $\rho$. By treating both $\mathbf{A}_{(1)}$ and $\text{Ker}\rho$ as linear algebra, one can obtain the basis of $\mathbf{A}_{(1)}/\text{Ker}\rho$ straightfowardly. This corresponds to the hierarchical structure of a topologically ordered state by applying CFT/TQFT correspondence (FIG. 1). The procedure taking quotient is nothing but gapping anyons belonging to $\text{Ker}\rho$ or assuming ordering $s_{(1)}=0$ for all the element $s_{(1)}\in\text{Ker}\rho$. Because of this operation, the algebraic relation is deformed to that of the IR theory and the condition $s_{(1)}=0$ is the anyon or mass condensation. Moreover, by studying the massless RGs successively, one can obtain the sequence of subalgebra structure and corresponding generalized symmetry breaking pattern. In this discussion, the construction of the ring homomorphism from $\mathbf{A}_{(1)}$ is fundamental and this can be constructed from the ideal decomposition of the ring $\mathbf{A}_{(1)}$. This argument can be thought of as a natural continuation of the research direction in [118, 119, 113].
Figure 1: Massless renormalization group and its connection to hierarchical structures in TOs. Under the CFT/TQFT, the homomorphism $\rho:\mathbf{A}_{(1)}\rightarrow\mathbf{A}_{(2)}$ induces hierarchical structure of anyons in TOs. Because of the existence of kernel of the homomorphism notified as $\text{Ker}\rho$ (which is intrinsically noninvertible), the fusion rule of a reduced theory $\mathbf{A}_{(2)}$ can become different from the original theory $\mathbf{A}_{(1)}$ and this can be understood as an emergent phenomenon[120](for more recent example, see the discussions and literature in [121, 122, 123]). In other words, $\mathbf{A}_{(2)}$ can be different from a subring of $\mathbf{A}_{(1)}$. This figure can be seen as an algebraic representation of the hierarchical structure in fractional quantum Hall states[124, 125, 126] and its connection to microscopic descriptions have been studied recently in [127, 128].
@@
where $Q$ is a symmetry operator (or integral of motion) representing the symmetry action of $\mathbf{A}_{(1)}$, i.e. $Q_{s^{\prime}_{(1)}}Q_{s_{(1)}}=Q_{s^{\prime}_{(1)}\times s_{(1)}}$. This means that the Hamiltonian and the symmetry operators (or integral of motions) are diagonalizable at the same time. Moreover, if $s_{(1)}$ is invertible, the ideal $\mathbf{I}_{(1)}$ becomes trivial because this means $\mathbf{A}_{(1)}=\mathbf{I}_{(1)}$. Hence, all the element $s_{(1)}\in\mathbf{I}_{(1)}$ is noninvertible. By interpreting the operator as a matrix[129], one can demonstrate that $Q_{s_{(1)}}$ has the eigenvalue 00. Hence by projecting the theory with $Q_{s_{(1)}}=0$, one can obtain the Hamiltonian $H_{(2)}$ with $\mathbf{A}_{(1)}/\mathbf{I}_{(1)}=\mathbf{A}_{(2)}$ symmetry. We note several references to study such matrix representation of (topological) symmetry operators in lattice models [130, 131, 132, 133, 134, 135, 136, 137, 138]. More practically, the careful treatment of emanant symmetry which distinguish the symmetry in lattice models and quantum field theoretic models (QFTs) will become important[139, 140, 141]. For the later discussion, we stress that an abelian algebra $\mathbf{A}_{(1)}$ can have a nontrivial (or nonabelian) ideal because one treats $\mathbf{A}_{(1)}$ as a group ring in mathematical terminology. In other words, the notion of abelian or invertibility depends on the detailed settings and representations.
@@
Based on the arguments above, we propose a systematic construction of systems with emergent symmetry (possibly noninvertible) from known existing lattice and QFT models (FIG. 2). This will be applicable to coupled wire systems[142, 143, 144, 145, 146, 147, 148, 149, 150, 151] or modulated symmetric systems[152, 115, 116, 116], for example. The ideal decomposition of abelian theory with a sufficient degree of freedom will be useful for the construction or classifications for example. Let us assume a Hamiltonian $H_{(1)}$ with the knowledge of symmetries $\mathbf{A}_{(1)}=\{\mathbf{A}_{(1)}^{[i]}\}_{i=1}^{N}$ and its ideal $\mathbf{I}_{(1)}$. Then, one can obtain the Hamiltonian $H_{(2)}$ with symmetry $\mathbf{A}_{(1)}/\mathbf{I}_{(1)}$ by representing the quantietion as a projection. For example, in a coupled wire model, the original model can be just a ladder model without inter-ladder interaction where the symmetry $\mathbf{A}_{(1)}^{[i]}$ acts on each ladder. However, by considering a nontrivial projection sending the ideal part to 00, one can obtain a more nontrivial model with inter-ladder interaction, and the symmetry (or integral of motions) is transformed to $\mathbf{A}_{(2)}$. For a modulated symmetric system, an analogous argument based on gauging (or orbifolding) can be seen in [153, 154], and corresponding arguments based on ideal will also be possible in principle.
@@
Figure 2: Construction flow of emergent or enhanced symmetries from coupled symmetries. From the knowledge of existing theories $\mathbf{A}_{(1)}^{[i]}$ and the ideal of the coupled symmetry $\mathbf{A}_{(1)}$, one can obtain system with nontrivial or interacting symmetry $\mathbf{A}_{(2)}$. Only by assuming the existence of the commutativity of the kernel, one can realize the algebraic operation as projection. Applications of this method to extended symmetry such as Haagerup symmetry or magma in E-series models [155] are interesting for future research. We note some research on the general construction of ideals in polynomial rings which might be useful for the general construction of ideals[156, 157].
@@
For the application of our method to the classification of the gapped phases, we remind that the massive RGs have a close connection to boundary conformal field theories (BCFTs)[158, 159, 160, 161, 162]. The correspondence has been studied in a wide variety of communities in various names as Moore-Seiberg data[163, 164, 165, 166], Li-Haldane conjecture[167, 160], and boson condensation[168, 22]. The combination of our algebraic arguments and symmetry analysis of boundary state in [169] will provide the general criteria to describe the general class of gapped phases. Related arguments can be seen in [170, 65, 66, 171, 172, 173, 174]. However, as the author has clarified in [69], the label of boundary state can be outside of the original symmetry $\mathbf{A}_{(1)}$ or $\mathbf{A}_{(2)}$. The corresponding boundary states are known as new boundary conditions in condensed matter[175, 176, 177] and symmetry-breaking boundary conditions in high energy or mathematical physics[178, 179, 180] because these boundary condition breaks extended bulk symmetry such as $W$ or Lie group symmetry in the bulk CFT (we note [181] as a useful reference).

### v2

@@ Abstract
We propose a general quantum Hamiltonian formalism of a renormalization group (RG) flow with an emphasis on generalized symmetry by interpreting the elementary relationship between homomorphism, quotient ring, and projection. In our formalism, the noninvertible nature of the ideal of a fusion ring realizing the generalized symmetry of an ultraviolet (UV) theory plays a fundamental role in determining condensation rules between anyons, resulting in the infrared (IR) theories. Our algebraic method applies to the domain wall problem in $2+1$ dimensional topologically ordered systems and the corresponding classification of $1+1$ dimensional gapped phase, for example. An ideal decomposition of a fusion ring provides a straightforward but strong constraint on the gapped phase with noninvertible symmetry and its symmetry-breaking (or emergent symmetry) patterns. Moreover, even in several specific homomorphisms connected under massless RG flows, less familiar homomorphisms appear, and we conjecture that they correspond to partially solvable models in recent literature. Our work demonstrates the fundamental significance of the abstract algebraic structure, ideal, for the RG in physics.
@@
Classification of the topological orders (TOs) and their fusion rules (or fusion rings[1]) between anyons is one of the most fundamental interests in contemporary condensed matter physics and related research fields[2, 3, 4, 5, 6]. Recently, the transformation law of the anyons in topologically ordered systems through a domain wall, which has been studied in the Kondo or impurity problem[7, 8, 9, 10, 11, 12, 13, 14, 15] traditionally, has captured attention in the fields as a potential classification framework of TOs more commonly (We note [16, 17, 18] as useful references for traditional research directions. We also note older works in astrophysics[19, 20] from a historical view). There exist mutually related ideas to study such settings, such as gapped or symmetry-preserving domain walls[21, 22, 23, 24, 25, 26, 27], Witt equivalence[28, 29, 30], and renormalization group (RG) domain walls[31, 32]. More practically, it is known that the theory containing a domain wall can explain the controversy of thermal Hall conductance in experimental settings[33, 34] and this is a consequence of the anomaly-inflow mechanism[20, 35, 36, 37]. In other words, if anyons in one theory can smoothly transform into anyons in another theory, these two theories are connected in some sense, and a nontrivial domain wall between them can exist. In this sense, bulk edge correspondence[2, 5] or the correspondence between conformal field theory (CFT) and topological quantum field theory (TQFT)[38], CFT/TQFT in short, should be relaxed to this connectivity or equivalence relations when applying the field-theoretic ideas to a realistic system as in experimental settings.
@@
On the other hand, such connectivity between theories is a close connection to symmetry-protected topological order (SPT) [39, 40, 41, 42, 43, 44, 45, 46, 47, 47, 48, 49, 50, 51, 52, 53] and its gapless analog called gapless SPT[54, 55, 56, 57] by the CFT/TQFT correspondence[38]. It has been gradually established that t’Hooft anomaly[58] or Lieb-Shultz-Mattis anomaly[59, 60] classification governs such connectivity between theories[61, 62, 63, 64, 65, 66, 67, 68] and the corresponding phenomena has been studied numerically in [69, 70, 71]. The most fundamental object in this classification is the integer spin simple current proposed by Schellekens and the collaborators[72, 73, 74, 75, 76, 77, 78, 79] which initially appeared in the study of discrete torsion[80, 81]. If the product of two theories forms an integer (or half-integer) spin simple current analogous to a hadron, the product theory can be extended or gauged with respect to this nontrivial integer spin simple current. By applying the folding trick[11], the product or coupled theory generates a class of domain walls between the theories phenomenologically. Hence, one may expect a systematic construction of anyon transformation laws between theories in the same anomaly class under such anomaly classification by the integer spin simple current. However, the research has still been limited at this stage (Probably, the fundamental problem in this research direction is the appearance of mixing of chirality of anyons in the coupled theories. To our knowledge, this object has been studied in the context of the fractional quantum Hall effect in higher Landau levels or anti-Pfaffian states [82, 83, 84]. See also the discussions in [85, 68]).
@@
In this work, we demonstrate a general phenomenology realizing an RG as a projection. For the construction of this projection, a mathematical structure called ideal plays a fundamental role. An ideal is fundamentally noninvertible and nonabelian (or nongroup-like). Moreover, it is also different from a subring, which commonly appears in physics literature. In principle, one can construct (infinitely) many new transformation laws of anyons by considering the ideal decomposition of existing fusion rings. More precisely, we study anyon transformation law or ring homomorphism between anyons in conformal field theories connected by massless RG flows[86]. Surprisingly, even in the simplest flow, such as the $SU(2)_{1}\times SU(2)_{1}$ Wess-Zumino-Witten (WZW) model to the $SU(2)_{2}$ WZW model and the flow from the tricritical Ising CFT to the Ising CFT, negative and fractional fusion coefficient, and unusual (or nonintegrable) flows appear inevitably. The unusual fusion coefficient has appeared recently in [87, 88, 57, 89, 90], and the unusual flow seems to correspond to partially solvable systems in [91, 92, 93].
@@
The rest of the manuscript is organized as follows. In section II, we provide a general or mathematical understanding of massless RG flow and its implications for gapped phases in $1+1$ dimensions and for bulk states in topological orders in $2+1$ dimensions. This is the main section of this work, and we demonstrate the fundamental importance of further studies on the ideal structure in the fusion ring. The argument is purely algebraic, and one can generalize the argument to higher-dimensional systems. In section III, we introduce a way of constructing a series of ideals from a given ring corresponding to a CFT. The method itself is well known in mathematical literature, but its implications for theoretical physics are quite strong. This construction itself implies the existence of (infinitely) many new series of massless RGs from a given CFT. In section IV, we provide several detailed constructions of homomorphisms between fusion rings appearing in CFTs connected under massless RG flows. We demonstrate the existence of nontrivial homomorphisms that are outside of existing integrable massless RG flow and propose its possible connection to partially solvable systems in recent literature[91, 92, 93]. Section. V is the concluding remarks of this manuscript, and we remark on a possible extension of our research and related phenomena that will be realizable in experimental settings. In Appendix A, we present an alternative way to obtain a Hamiltonian realizing the massless RG by taking some limit. The resultant model can be interpreted as a gauge model involving the ideal structures. In Appendix B, we provide a conjectural decomposition of CFTs related by massless RG, which will be useful in studying the homomorphism between models.
I.1 Summary of the main ideas
In this work, we revisit existing formalisms with a less familiar (but traditional) way and formulate several frameworks in a rigorous algebraic way. Before going into details, we summarize the main claims in this manuscript.
  • A massless RG in the quantum Hamiltonian formalism can be formulated as a projection at the level of linear algebra. (Eq. (4) and Fig.1.)
  • The structure of ideals of a given fusion ring (or UV theory) produces a general method for generating massless RG and the resultant IR theories. (Fig. 4)
  • We provide a rigorous interpretation of sandwich construction in literature as implications of the massless RG to the massive RG. (Fig. 3 and Eq. (9))
  • Only by assuming the UV and IR fusion rings, the homomorphism is not unique.
We stress that the construction of IR or emergent theory can be determined only by assuming the UV theory, and the systematic construction of such an IR theory has not been achieved in the literature. By applying the CFT/TQFT correspondence, the above points provide a classification of fusion rings of anyons, or SymTFTs in TOs. There exist several theoretical studies assuming the homomorphism (or functor) in the communities (See literature in [85, 68]), but our work is the first general framework for constructing such homomorphisms.
II Substructure from homomorphism: Projection as renormalization group flow
@@
where $\mathbf{A}$ is an algebra describing generalized symmetries (or integral of motions) or underlying fusion category of anyons, and the lower index $(1)$ and $(2)$ label the ultra-violet (UV) and infra-red (IR) theories, respectively. The corresponding argument has appeared evidently in [32] based on the earlier work [95] and the folding trick [11]. When interpreting this homomorphism as a defect between two theories, it is called the RG domain wall, and it is expected to describe a massless RG[86]. We also note that the categorical structure corresponding to the ring homomorphism, called a functor, has been studied as gapped (or symmetry-preserving) domain walls in [22, 23, 25, 27] with emphasis on the theoretical aspect of TQFTs or TOs.
@@
We provide an interpretation of the algebraic analog of a coset of $\mathbf{A}_{(1)}$ by $\text{Ker}\rho$ or the quotient ring $\mathbf{A}_{(1)}/\text{Ker}\rho$ which can be constructed from the homomorphism $\rho$ corresponding to the massless RGs. We note that the idea of introducing a coset by group symmetry, known as orbifolding, has been studied widely[96, 97, 98, 80, 99, 100, 101] and is generalized to nonabelian symmetry[102, 103], but the quotient ring which we will discuss is different from them. The object, $\text{Ker}\rho$, dividing a ring is an ideal, and a nontrivial ideal only forms a pseudo ring or rng without identity. (We cite [104] as a recent English translation of the fundamental work by Noether [105] to emphasize the fundamental importance of ideals in contemporary mathematics.) In other words, an ideal is intrinsically noninvertible or nonabelian (or non-group-like). From these properties, it should be distinguished from a subring, which is also familiar in physics literature. We also note that the symbol “//" in this section is different from “//" in coset CFTs[106, 107, 108, 109], whereas there exist some connections between the models as in [110, 32] (We introduce a conjectural relationship between coset CFTs and massless RGs in the Appendix. B). As far as we know, the term ideal appeared in physics literature in [111, 1, 112, 113] as fusion ring ideal, in [114] in the study of matrix product operator, in [115] with a close connection to massless RG flows and in [116] in studying polynomial ring representations of fractonic systems[117, 118].
@@
We restrict our attention to abstract algebra, not to category theory, because the underlying category theories can be premodular fusion or superfusion categories, or even outside of them that are still under development. However, the basic fusion ring in this manuscript is well-defined and has been studied several times both in the mathematics and physics communities. In other words, our discussion can be a clue to constructing a new fundamental class of categories. The fusion rings with negative or fractional coefficients are the corresponding structure, and they appeared in the study of fermionic string theories [119] and more recently in [88, 57], for example. The corresponding fusion rings have been studied more generally by the first author in [89, 90] with a close connection to the simple current extension by Schellekens and the collaborators[72, 73, 74, 75, 76, 77, 78, 79]. We also note that the topological symmetry operator (or integral of motion) is an operator over complex fields. Hence, the quantum Hamiltonian formalism of the symmetry operator itself does not restrict the coefficients to nonnegative integer coefficients of the fusion rings. The appearance of a nonnegative integer coefficient is the consequence of the correspondence between the symmetry operator and the defects under the topological Wick rotation (or modular $S$ transformation in two-dimensional CFTs)[131] [^1] We thank Yuji Tachikawa, Jurgen Fuchs, and Shinichiro Yahagi for the related discussions..
@@
where $\text{Ker}\rho$, forming an ideal, is a kernel of the map $\rho$ and the inclusion is defined by interpreting the objects as linear algebra (not as a ring). This is widely known as the ring homomorphism theorem in mathematics (or an example of Noether’s isomorphism theorems), and one can see the profound history initiated by Kummer, Dedekind, and more. In the present theoretical physics context, this means the correspondence between $\mathbf{A}_{(1)}/\text{Ker}\rho$ and the IR generalized symmetries (or integral of motions) of massless flow $\rho$. By treating both $\mathbf{A}_{(1)}$ and $\text{Ker}\rho$ as linear algebra, one can obtain the basis of $\mathbf{A}_{(1)}/\text{Ker}\rho$ straightfowardly. This corresponds to the hierarchical structure of a topologically ordered state by applying CFT/TQFT correspondence (FIG. 1). The procedure taking the quotient is nothing but gapping anyons belonging to $\text{Ker}\rho$ or assuming ordering $s_{(1)}=0$ for all the elements $s_{(1)}\in\text{Ker}\rho$. Because of this operation, the algebraic relation is deformed to that of the IR theory, and the condition $s_{(1)}=0$ is the anyon or mass condensation. Moreover, by studying the massless RGs successively, one can obtain the sequence of subalgebra structures and the corresponding generalized symmetry-breaking pattern. In this discussion, the construction of the ring homomorphism from $\mathbf{A}_{(1)}$ is fundamental, and this can be constructed from the ideal decomposition of the ring $\mathbf{A}_{(1)}$. This argument can be thought of as a natural continuation of the research direction in [120, 121, 115].
Figure 1: Massless renormalization group and its connection to hierarchical structures in TOs. Under the CFT/TQFT, the homomorphism $\rho:\mathbf{A}_{(1)}\rightarrow\mathbf{A}_{(2)}$ induces hierarchical structure of anyons in TOs. Because of the existence of kernel of the homomorphism notified as $\text{Ker}\rho$ (which is intrinsically noninvertible), the fusion rule of a reduced theory $\mathbf{A}_{(2)}$ can become different from the original theory $\mathbf{A}_{(1)}$ and this can be understood as an emergent phenomenon[122](for more recent example, see the discussions and literature in [123, 124, 125]). In other words, $\mathbf{A}_{(2)}$ can be different from a subring of $\mathbf{A}_{(1)}$. This figure can be seen as an algebraic representation of the hierarchical structure in fractional quantum Hall states[126, 127, 128] and its connection to microscopic descriptions has been studied recently in [129, 130].
@@
where $Q$ is a symmetry operator (or integral of motion) representing the symmetry action of $\mathbf{A}_{(1)}$, i.e. $Q_{s^{\prime}_{(1)}}Q_{s_{(1)}}=Q_{s^{\prime}_{(1)}\times s_{(1)}}$. This means that the Hamiltonian and the symmetry operators (or integral of motions) are diagonalizable at the same time. Moreover, if $s_{(1)}$ is invertible, the ideal $\mathbf{I}_{(1)}$ becomes trivial because this means $\mathbf{A}_{(1)}=\mathbf{I}_{(1)}$. Hence, all the element $s_{(1)}\in\mathbf{I}_{(1)}$ is noninvertible. By interpreting the operator as a matrix[131], one can demonstrate that $Q_{s_{(1)}}$ has the eigenvalue 0. Hence by projecting the theory with $Q_{s_{(1)}}=0$, one can obtain the Hamiltonian $H_{(2)}$ with $\mathbf{A}_{(1)}/\mathbf{I}_{(1)}=\mathbf{A}_{(2)}$ symmetry. We note several references to study such matrix representation of (topological) symmetry operators in lattice models [132, 133, 134, 135, 136, 137, 138, 139, 140]. More practically, the careful treatment of emanant symmetry which distinguish the symmetry in lattice models and quantum field theoretic models (QFTs) will become important[141, 142, 143]. For the later discussion, we stress that an abelian algebra $\mathbf{A}_{(1)}$ can have a nontrivial (or nonabelian) ideal because one treats $\mathbf{A}_{(1)}$ as a group ring in mathematical terminology. In other words, the notion of abelian or invertibility depends on the detailed settings and representations.
@@
Based on the arguments above, we propose a systematic construction of systems with emergent symmetry (possibly noninvertible) from known existing lattice and QFT models (FIG. 2). This will be applicable to coupled wire systems[144, 145, 146, 147, 148, 149, 150, 151, 152, 153] or modulated symmetric systems[154, 117, 118, 118], for example. The ideal decomposition of an abelian theory with a sufficient degree of freedom will be useful for the construction or classification, for example. Let us assume a Hamiltonian $H_{(1)}$ with the knowledge of symmetries $\mathbf{A}_{(1)}=\{\mathbf{A}_{(1)}^{[i]}\}_{i=1}^{N}$ and its ideal $\mathbf{I}_{(1)}$. Then, one can obtain the Hamiltonian $H_{(2)}$ with symmetry $\mathbf{A}_{(1)}/\mathbf{I}_{(1)}$ by representing the quotient as a projection. For example, in a coupled wire model, the original model can be just a ladder model without inter-ladder interaction where the symmetry $\mathbf{A}_{(1)}^{[i]}$ acts on each ladder. However, by considering a nontrivial projection sending the ideal part to 0, one can obtain a more nontrivial model with inter-ladder interaction, and the symmetry (or integral of motions) is transformed to $\mathbf{A}_{(2)}$. For a modulated symmetric system, an analogous argument based on gauging (or orbifolding) can be seen in [155, 156], and corresponding arguments based on ideal will also be possible in principle.
@@
Figure 2: Construction flow of emergent or enhanced symmetries from coupled symmetries. From the knowledge of existing theories $\mathbf{A}_{(1)}^{[i]}$ and the ideal of the coupled symmetry $\mathbf{A}_{(1)}$, one can obtain system with nontrivial or interacting symmetry $\mathbf{A}_{(2)}$. Only by assuming the existence of the commutativity of the kernel, one can realize the algebraic operation as a projection. Applications of this method to extended symmetry, such as Haagerup symmetry or magma in E-series models [157] are interesting for future research. We note some research on the general construction of ideals in polynomial rings, which might be useful for the general construction of ideals[158, 159].
@@
For the application of our method to the classification of the gapped phases, we remind that the massive RGs have a close connection to boundary conformal field theories (BCFTs)[160, 161, 162, 163, 164]. The correspondence has been studied in a wide variety of communities under various names as Moore-Seiberg data[165, 166, 167, 168], Li-Haldane conjecture[169, 162], and boson condensation[170, 21]. The combination of our algebraic arguments and symmetry analysis of boundary state in [171] will provide the general criteria to describe the general class of gapped phases. Related arguments can be seen in [172, 64, 65, 173, 174, 175, 176]. However, as the author has clarified in [68], the label of boundary state can be outside of the original symmetry $\mathbf{A}_{(1)}$ or $\mathbf{A}_{(2)}$. The corresponding boundary states are known as new boundary conditions in condensed matter[177, 178, 179] and symmetry-breaking boundary conditions in high energy or mathematical physics[180, 181, 182] because these boundary condition breaks extended bulk symmetry such as $W$ or Lie group symmetry in the bulk CFT (we note [183] as a useful reference). For the later convenience, we introduce the extended ring (or extended algebra) $\mathbf{A}^{\text{ex}}$ with $\mathbf{A}\subset\mathbf{A}^{\text{ex}}$ and assume that the underlying CFT Hamiltonians are symmetric under $\mathbf{A}^{\text{ex}}$ action but the Hilbert spaces are labeled by $\mathbf{A}^{\text{ex}}$. Then one can express the set $\mathbf{R}\subset\mathbf{A}^{\text{ex}}$ labelling $\mathbf{A^{\prime}}$ invariant module as follows,
[…]
Hence, in this formalism, one can formulate $\mathbf{R}$ as a generalization of an ideal because $r\in\mathbf{R}$ can be outside of $\mathbf{A^{\prime}}$. We propose that the above subring invariant set governs the massive RG and their classifications. We remind that whereas the set is invariant under the action of the subring, the object will be noninvariant, and this is nothing but the SSB.